Shaun and Tom leave their house at the same time riding their bicycles in opposite directions. Shaun rides four mph faster than Tom. After three hours, they are 64.2 miles apart. At what speed was each person riding?
Answer by lwsshak3(11628)

Shaun and Tom leave their house at the same time riding their bicycles in opposite directions. Shaun rides four mph faster than Tom. After three hours, they are 64.2 miles apart. At what speed was each person riding?
let x=speed at which tom is traveling
x+4=speed at which Shaun is traveling
speed at which they are traveling away from each other=sum of their indv speeds, x+x+4=2x+4
distance/speed=travel time
64.2/(2x+4)=3
6x+12=64.2
6x=52.2
x=8.7
x+4=12.7
speed at which tom is traveling=8.7 mph
speed at which Shaun is traveling=12.7 mph